OpenStudy (anonymous):

I am still looking for some help please. I am completely confused.. Find f-1(x) (THE -1 IS AN INVERSE FUNCTION). Sketch the graphs of y=f(x) and y=f-1(x) on the same coordinate axes. Answer for f(x)=3x-4 I am not understanding how this works...

OpenStudy (anonymous):

for the inverse switch x and y so your original formula is y = 3x - 4 now switch x and y for inverse x = 3y - 4 this is a correct answer but teachers usually want it in y = mx + b form so y = (x + 4)/3 would be your answer

OpenStudy (anonymous):

What if I tried f(x)=x^3+1

OpenStudy (anonymous):

How are exponents carried over in these types of equations?

OpenStudy (anonymous):

Isolate it one side. The inverse of cube is the cube root

OpenStudy (anonymous):

To sketch the graph take the values of x = -3 to x = 3 and substitute into the equation f(x) = 3x +4 and plot the point therefore when x = -3 f(x) or y = -4 when x= -2 f(x) or y = -2 when x =-1 f(x) or y = 1 when x = 0 y = 4 when x = 1 y = 7 when x = 2 y =10 when x = 3 y = 11 plot those points then do the same for the inverse function .